# Quick on the Uptake: Eliciting Implicit Intents from Human Demonstrations for Personalized Mobile-Use Agents

### v1

@@ Abstract
As multimodal large language models advance rapidly, the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions from graphical user interface. To further enhance mobile-use agents, previous studies employ demonstration learning to improve mobile-use agents from human demonstrations. However, these methods focus solely on the explicit intention flows of humans (e.g., step sequences) while neglecting implicit intention flows (e.g., personal preferences), which makes it difficult to construct personalized mobile-use agents. In this work, to evaluate the Intention Alignment Rate between mobile-use agents and humans, we first collect MobileIAR, a dataset containing human-intent-aligned actions and ground-truth actions. This enables a comprehensive assessment of the agents’ understanding of human intent. Then we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. IFRAgent then leverages a SOP extractor combined with retrieval-augmented generation and a query rewriter to generate personalized query and SOP from a raw ambiguous query, enhancing the alignment between mobile-use agents and human intent. Experimental results demonstrate that IFRAgent outperforms baselines by an average of 6.79% (32.06% relative improvement) in human intention alignment rate and improves step completion rates by an average of 5.30% (26.34% relative improvement). The codes are available at https://github.com/MadeAgents/Quick-on-the-Uptake.
Introduction
@@
Figure 1: Comparing IFRAgent with normal mobile-use agents and existing demonstration learning methods for mobile-use agents. IFRAgent considers both explicit intention flow and implicit intention flow, enabling it to capture more personalized information, such as taste preferences.
@@
As multimodal large language models advance rapidly (Zhang et al. 2024b; Yin et al. 2024), the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions (e.g., clicking, scrolling) from graphical user interface (GUI) (Zhang et al. 2024a; Liu et al. 2025b). To further improve capabilities of mobile-use agents, some works (Liu et al. 2025a; Verma et al. 2024) employ demonstration learning to enable mobile-use agents to know how to complete task from human demonstrations.
@@
However, as show in Figure 1, existing demoinstration learning method for mobile-use agents focus solely explicit intention flows of humans (e.g., operational logic, step sequences) to help mobile-use agents learn how to complete task. Moreover, user instructions in the real world are often ambiguous (Cheng et al. 2025b) and user-specific, requiring mobile-use agents to understand the implicit intention flows of humans in order to align with human intentions.
@@
For challenge (i), we first collect MobileIAR, a user-specific dataset that supports both English and Chinese language users, designed to assess the Intention Alignment Rate between mobile-use agents and humans. As shown in Table 1, the dataset contains 16 apps and 945 instuction spanning 7 daily scenarios. And the dataset provides both the human-intent-aligned actions and ground-truth action lists, enabling a comprehensive assessment of the alignment level between mobile-use agents and human intentions.
For challenge (ii), we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. The IFRAgent consists of the intention flow extraction phase and the deployment phase. In the intention flow extraction phase, IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. In the deployment phase, IFRAgent, as a plug-and-play module, leverages a SOP extractor combined with retrieval-augmented generation (RAG) (Lewis et al. 2020) and a warmed-up query rewriter to rewrite the user’s ambiguous query into a user-specific personalized query and a personalized SOP based on the analysis of the human intention flow from the previous phase, thereby enabling mobile-use agents to align with human intentions.
Extensive experiments spanning diverse mobile-use agents (supporting multiple user languages and are constructed with varying methods) demonstrate that IFRAgent outperforms baseline methods by an average of 6.79% (32.06% relative improvement) in intent alignment rate and achieves a 5.30% in task completion rate (26.34% relative improvement). We also find that general-domain models (e.g., Qwen2.5-VL-7B, GPT-4o) demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS.
We further validate IFRAgent’s capability and generalizability through ablation study, cross-dataset tests, comparative studies with other methods, and scale analysis.
@@
(iii) Through extensive experiments on different mobile-use agents, we demonstrate that IFRAgent achieves an average improvement of 6.79% (32.06% relative improvement) in intent alignment rate and 5.30% in step-wise success rate (26.34% relative improvement) over baseline methods.
@@
IFRAgent 945 16 7.7
Table 1: Comparative analysis of datasets and environments for evaluating mobile-use agents. Key metrics include: # Inst. (instruction count), # Apps (application count), # Step (average steps per task), HL (high-level instructions), GT (ground truth trajectories), FS (few-shot learning support), and US (user-specific demonstrations). Data for this table is partially sourced from Liu et al. (2025a).
Related work
Mobile-use agents for mobile automation
@@
Mobile-use agents (Wang et al. 2024b; Hu et al. 2025) can automatically perform mobile tasks by using the GUI to simulate human interactions (e.g., clicking, scrolling). Researchers mainly adopt two approaches to build mobile-use agents: using open-source models or using closed-source models. Some construct mobile-use (M)LLMs with open-source models  (Zhang and Zhang 2024; Hong et al. 2024; Wu et al. 2025; Xu et al. 2024b; Qin et al. 2025), often followed by post-training via reinforcement learning  (Wang et al. 2024c; Zhou et al. 2024; Luo et al. 2025; Gu et al. 2025). Others rely on the general-domain knowledge of closed-source models to develop mobile-use agents  (Jiang et al. 2025; Wang et al. 2025; Li et al. 2025). Both approaches have demonstrated strong performance in mobile automation tasks. As shown in Table 1, there are many benchmarks for evaluating mobile-use agents. These benchmarks mainly fall into two types: those that use static images for testing (Shaw et al. 2023; Li et al. 2024; Rawles et al. 2023; Zhang et al. 2024c) and those that employ dynamic environments to evaluate (Wang et al. 2024a; Zhang et al. 2025; Rawles et al. 2025; Xu et al. 2024a). However, these benchmarks only assess whether mobile-use agents complete tasks, lacking a dataset that can assess the alignment level between mobile-use agents and human intentions.
Figure 2: The workflow pipeline of IFRAgent. Above the dashed line shows the intention flow extraction phase. Below the dashed line shows the deployment phase.
Demonstration learning for mobile-use agents
@@
Demonstration learning (Correia and Alexandre 2024) is a method that enhances the capabilities of agents by observing human demonstrations, primarily including imitation learning (Rybski et al. 2007) and inverse reinforcement learning (Ng, Russell et al. 2000). This approach has been widely applied in the robot domain (Argall et al. 2009), and in the field of mobile-use and computer-use agents, some works have utilized demonstration learning to improve agents’ ability to accomplish tasks through few-shot examples (Liu et al. 2025a; Verma et al. 2024). However, existing demonstration learning research in the mobile-use agent domain has focused solely on the explicit intention flow in human demonstrations while neglecting implicit intention flows such as user behavioral preferences. IFRAgent comprehensively analyzes both the explicit and implicit intention flows in human demonstrations, thereby establishing a more user-intent-aligned paradigm for mobile-use agent demonstration learning.
IFRAgent Framework
@@
We propose IFRAgent, a framework that enhances intent alignment between mobile-use agents and humans. Figure 2 shows an overview. In this section, we first introduce the pipeline framework and then introduce the trainable scheme.
IFRAgent pipline
@@
The IFRAgent pipeline can be divided into 2 phases: the intention flow extraction phase and the deployment phase. Their algorithmic representations are shown in Appendix.
Intention Flow Extraction Phase
@@
In the intention flow extraction phase, IFRAgent collects human demonstrations and analyze these human demonstrations to extract the implicit intention flow and explicit intention flow of humans. For each user $u_{i}\in U=\{u_{1},u_{2},\dots,u_{n}\}$, we first collect human demonstrations comprising a set of queries $Q_{i}=\{q_{1},q_{2},\dots,q_{k}\}$ and initialize an empty user-level habit repository $h_{i}$. Each query $q_{j}\in Q_{i}$ is accompanied by a sequence of operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\dots,s_{p}\}$ provided by the user.
@@
Deployment Phase
$(q^{\prime},p^{\prime})=\begin{cases}(q_{j},p_{j})&\text{if }\exists j=\arg\max_{k}\text{sim}(\mathbf{l},\mathbf{l}_{k})\\ &\quad\land\ \text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau,\\ \varnothing&\text{otherwise.}\end{cases}$ (3)
$p=\mathcal{E}(q,(q^{\prime},p^{\prime})).$ (4)
$(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i}).$ (5)
$a=\mathcal{F}(\hat{q},\hat{p},s).$ (6)
Trainable Scheme
The SOP Extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are two key components of the IFRAgent during the deployment phase. Both $\mathcal{E}$ and $\mathcal{W}$ are models deployable on the edge side, thus lacking general knowledge about mobile operations. Since $\mathcal{E}$ has already compensated for the lack of general knowledge through few-shot learning with RAG, so only $\mathcal{W}$ require supervised fine-tuning to unleash its potential in query rewriting.
To warm up the trainable scheme query writer $\mathcal{W}$, we first employ a crowdsourcing approach to directly collect data from a group of English and Chinese users, who fill out the habit repository $H=\{h_{1},h_{2},\dots,h_{n}\}$. Then, using a combination of manual construction and LLM-generated expansions, we create a set of ambiguous instructions $q$. These users are asked to provide their own SOP $p$ for each ambiguous instruction, as well as their personalized rewritten queries $\hat{q}$ and personalized rewritten SOPs $\hat{p}$.
Next, we train the query writer $\mathcal{W}$ to predict the personalized rewritten query $\hat{q}$ and the personalized rewritten SOP $\hat{p}$, given the user $u_{i}$’s habit repository $h_{i}$, the ambiguous instruction $q$, and the corresponding SOP $p$. We use this dataset to train the query writer $\mathcal{W}$ for one epoch to achieve the warm-up effect.
The training objective can be formulated as:
[…]
After warm-up training, $\mathcal{W}$ acquires the capability to rewrite queries and SOPs in a user-specific manner.
Experiments
@@
                    +IFRAagent 20.370.51↑ 81.821.91↑ 17.560.87↑ 24.042.04↑ 84.422.87↑ 21.342.30↑
Table 2: Performance comparison of mobile-use agents with IFRAagent enhancements, categorized by open-source and close-source models. IFRAgent demonstrates improvements across nearly all metrics.
Experiments Setup
Baseline

### v2

@@ Abstract.
As multimodal large language models advance rapidly, the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions from graphical user interfaces. To further enhance mobile-use agents, previous studies employ demonstration learning to improve mobile-use agents from human demonstrations. However, these methods focus solely on the explicit intention flows of humans (e.g., step sequences) while neglecting implicit intention flows (e.g., personal preferences), which makes it difficult to construct personalized mobile-use agents. In this work, to evaluate the Intention Alignment Rate between mobile-use agents and humans, we first collect MobileIAR, a dataset containing human-intent-aligned actions and ground-truth actions. This enables a comprehensive assessment of the agents’ understanding of human intent. Then we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. IFRAgent then leverages a SOP extractor combined with retrieval-augmented generation and a query rewriter trained through knowledge distillation to generate personalized queries and SOPs from a raw ambiguous query, enhancing the alignment between mobile-use agents and human intent. Experimental results demonstrate that IFRAgent outperforms baselines by an average of 28.80% relative improvement in human intention alignment rate and improves step completion rates by an average of 24.20% relative improvement.
GUI agent, MLLM agent, Personalized agent, Mobile-use agent
submissionid: 417copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Figure 1. Comparing IFRAgent with normal mobile-use agents and existing demonstration learning methods for mobile-use agents. IFRAgent considers both explicit intention flow and implicit intention flow, enabling it to capture more personalized information, such as taste preferences.
@@
As multimodal large language models advance rapidly (Zhang et al., 2024b; Yin et al., 2024; Chen et al., 2026), the automation of mobile tasks has become increasingly feasible through the use of mobile-use agents that mimic human interactions (e.g., clicking, scrolling) from graphical user interface (GUI) (Zhang et al., 2024a; Liu et al., 2025b; Tang et al., 2025b). To further improve the capabilities of mobile-use agents, some works (Liu et al., 2025a; Verma et al., 2024) employ demonstration learning to enable mobile-use agents to know how to complete tasks from human demonstrations.
@@
However, as shown in Figure 1, existing demonstration learning methods for mobile-use agents focus solely on explicit human intention flows (e.g., operational logic, step sequences) to help mobile-use agents learn how to complete tasks. Moreover, user instructions in the real world are often ambiguous (Cheng et al., 2025b) and user-specific, requiring mobile-use agents to understand the implicit intention flows of humans in order to align with human intentions.
@@
For challenge (i), we first collect MobileIAR. This user-specific dataset supports both English and Chinese language users, designed to assess the Intention Alignment Rate between mobile-use agents and humans. The dataset contains 16 apps and 945 instructions spanning seven daily scenarios. Moreover, the dataset provides both the human-intent-aligned actions and ground-truth action lists, enabling a comprehensive assessment of the alignment level between mobile-use agents and human intentions.
For challenge (ii), we propose IFRAgent, a framework built upon Intention Flow Recognition from human demonstrations. The IFRAgent consists of the intention flow extraction phase and the deployment phase. In the intention flow extraction phase, IFRAgent analyzes explicit intention flows from human demonstrations to construct a query-level vector library of standard operating procedures (SOP), and analyzes implicit intention flows to build a user-level habit repository. In the deployment phase, IFRAgent, as a plug-and-play module, leverages a SOP extractor combined with retrieval-augmented generation (RAG) (Lewis et al., 2020) and a knowledge distillation warmed-up query rewriter to rewrite the user’s ambiguous query into a user-specific personalized query and a personalized SOP based on the analysis of the human intention flow from the previous phase. This enables mobile-use agents to better align with human intentions, while also improving their general task completion capability to a certain extent.
Extensive experiments spanning diverse mobile-use agents (supporting multiple user languages and constructed with varying methods) demonstrate that IFRAgent outperforms baseline methods by an average of 6.79% (28.80% relative improvement) in intent alignment rate and achieves a 5.30% in task completion rate (24.20% relative improvement). We also find that general-domain models (e.g., Qwen2.5-VL, GPT-4o) demonstrate more significant improvements compared to specialized mobile-use agent base models.
We further validate IFRAgent’s capability and generalizability through an ablation study, cross-dataset tests, comparative studies with other methods, and scale analysis.
@@
(iii) Through extensive experiments on different mobile-use agents, we demonstrate that IFRAgent achieves an average improvement of 6.56% (32.06% relative improvement) in intent alignment rate and 5.50% in step-wise success rate (26.34% relative improvement) over baseline methods.
Table 1. Comparative analysis of datasets and environments for evaluating mobile-use agents. Key metrics include: # Inst. (instruction count), # Apps (application count), # Step (average steps per task), HL (high-level instructions), GT (ground truth trajectories), FS (few-shot learning support), and US (user-specific demonstrations). Data for this table is partially sourced from Liu et al. (2025a).
@@
MobileIAR 945 16 7.7
Figure 2. Examples of human-intent-aligned actions and SR ground truth lists. Even for the same query, the human-intent-aligned actions of user 1, user 2, and user 3 are different. However, these actions, along with some other actions that are also beneficial for completing the query, all belong to the SR ground truth list.
Figure 3. Construction of the MobileIAR dataset. In Round 1, annotators independently label human intent-aligned actions. In Round 2, groups of three participants are formed to jointly label the SR (step-wise success rate) ground truth lists based on the trajectories obtained from Round 1.
2. Related work
2.1. Mobile-use agents for mobile automation
@@
Mobile-use agents (Wang et al., 2024b; Hu et al., 2025) can automatically perform mobile tasks by using the GUI to simulate human interactions (e.g., clicking, scrolling). Researchers mainly adopt two approaches to build mobile-use agents: using open-source models or using closed-source models. Some construct mobile-use (M)LLMs with open-source models  (Zhang and Zhang, 2024; Hong et al., 2024; Wu et al., 2025; Xu et al., 2024b; Qin et al., 2025), often followed by post-training via reinforcement learning  (Wang et al., 2024c; Zhou et al., 2024; Lu et al., 2026; Tang et al., 2025a). Others rely on the general-domain knowledge of closed-source models to develop mobile-use agents  (Jiang et al., 2025; Wang et al., 2025b; Li et al., 2025). Both approaches have demonstrated strong performance in mobile automation tasks. As shown in Table 1, there are many benchmarks for evaluating mobile-use agents. These benchmarks mainly fall into two types: those that use static images for testing (Shaw et al., 2023; Li et al., 2024; Rawles et al., 2023; Zhang et al., 2024c) and those that employ dynamic environments to evaluate (Wang et al., 2024a; Zhang et al., 2025; Rawles et al., 2025; Xu et al., 2024a). However, these benchmarks only assess whether mobile-use agents complete tasks, lacking a dataset that can assess the alignment level between mobile-use agents and human intentions.
2.2. Demonstration learning for mobile-use
@@
Demonstration learning (Correia and Alexandre, 2024) is a method that enhances the capabilities of agents by observing human demonstrations, primarily including imitation learning (Rybski et al., 2007) and inverse reinforcement learning (Ng et al., 2000). This approach has been widely applied in the robot domain (Argall et al., 2009), and in the field of mobile-use and computer-use agents, some works have utilized demonstration learning to improve agents’ ability to accomplish tasks through few-shot examples (Liu et al., 2025a; Verma et al., 2024). Another popular approach to building mobile-use agents via demonstration learning is to leverage videos (Zhang et al., 2026; Jung et al., 2026; Wang et al., 2025a). By efficiently processing videos, valuable data can be extracted to serve as references for mobile-use agents or to synthesize training data. However, existing demonstration learning research in the mobile-use agent domain has focused solely on the explicit intention flow in human demonstrations while neglecting implicit intention flows such as user behavioral preferences. IFRAgent comprehensively analyzes both the explicit and implicit intention flows in human demonstrations, thereby establishing a more user-intent-aligned paradigm for mobile-use agent demonstration learning.
3. MobileIAR dataset collection
In this section, we first explain why the emergence of MobileIAR is necessary, and then we introduce the specific process of constructing MobileIAR.
3.1. Motivation of MobileIAR dataset
In real-world applications, the same query given by different users can have different implicit meanings. Ideal mobile-use agents need to understand the implicit intent behind the user’s query, enabling personalized operation of the user’s smartphone. Such personalized mobile-use agents will significantly enhance the user experience and drive development in the agent-human interaction paradigm.
However, in order to construct personalized mobile-use agents, the first issue to address is how to assess the level of personalization and alignment between agents and human intentions. Currently, there is a lack of datasets or benchmarks that can evaluate the alignment between mobile-use agents and human intentions. Therefore, we propose the MobileIAR dataset and provide a measurement method for the IAR (Intention Alignment Rate) metric within MobileIAR, laying the foundation for personalized mobile-use agents. As shown in Figure 2, MobileIAR provides both human-intent-aligned actions to measure the degree of intent alignment, and ground truth lists to measure task completion capability.
3.2. Construction of MobileIAR dataset
To construct MobileIAR dataset, as shown in Figure 3, we first collect trajectories for both English-speaking and Chinese-speaking users across seven daily-life scenarios through a crowdsourcing approach, resulting in a total of 945 instructions and 7,310 screenshots. The annotators cover a diverse group of participants, including men and women from different age groups.
In the annotation process, we conduct two rounds of annotation. The first round of annotation is done independently by the annotators, where they collect trajectories on their own real mobile devices. We encourage the annotators to choose completion methods that highlight their personalized preferences for the same instruction. The results of this round are used as the human-intent-aligned action. The human-intent-aligned action measures the ability of the mobile-use agent to align with the user’s intent.
The second round of annotation builds upon the first round, aiming to allow MobileIAR to assess not only the mobile-use agent’s intent alignment capability but also its general competence. Since the human-intent-aligned action is not always the only correct action in some cases, and often the agent’s execution path may have multiple routes, we assign each trajectory from the first round randomly to three annotators to label possible correct alternative actions. The intersection of these labeled actions is then taken as the ground-truth action list. The ground-truth action list measures the mobile-use agent’s general task completion ability.
4. IFRAgent Framework
Figure 4. The workflow pipeline of IFRAgent. The phase above the dashed line represents the intention flow extraction phase, while the phase below the dashed line represents the deployment phase. In the intention flow extraction phase, IFRAgent extracts SOPs and user profiles based on users’ trajectories. In the deployment phase, IFRAgent uses this information to rewrite personalized queries and generate personalized SOPs.
@@ 4. IFRAgent Framework
We propose IFRAgent, a framework that enhances intent alignment between mobile-use agents and humans. Figure 4 shows an overview. In this section, we first introduce the pipeline framework and then introduce the trainable scheme via knowledge distillation.
4.1. IFRAgent pipline
@@
The IFRAgent pipeline can be divided into 2 phases: the intention flow extraction phase and the deployment phase.
4.1.1. Intention Flow Extraction Phase
@@
In the intention flow extraction phase, IFRAgent collects human demonstrations and analyzes these human demonstrations to extract the implicit intention flow and explicit intention flow of humans. For each user $u_{i}\in U=\{u_{1},u_{2},\dots,u_{n}\}$, we first collect human demonstrations comprising a set of queries $Q_{i}=\{q_{1},q_{2},\dots,q_{k}\}$ and initialize an empty user-level habit repository $h_{i}$. Each query $q_{j}\in Q_{i}$ is accompanied by a sequence of operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\dots,s_{p}\}$ provided by the user.
@@
Table 2. Performance comparison of mobile-use agents with IFRAagent enhancements, categorized by open-source and close-source models. IFRAgent demonstrates improvements across nearly all metrics.
@@
                    + IFRAgent 20.370.51↑ 81.821.91↑ 17.560.87↑ 24.042.04↑ 84.422.87↑ 21.342.30↑
4.1.2. Deployment Phase
(3) $(q^{\prime},p^{\prime})=\begin{cases}(q_{j},p_{j})&\text{if }\exists j=\arg\max_{k}\text{sim}(\mathbf{l},\mathbf{l}_{k})\\ &\quad\land\ \text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau,\\ \varnothing&\text{otherwise.}\end{cases}$
(4) $p=\mathcal{E}(q,(q^{\prime},p^{\prime})).$
(5) $(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i}).$
(6) $a=\mathcal{F}(\hat{q},\hat{p},s).$
4.2. Training via Knowledge Distillation
The SOP Extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are two key components of the IFRAgent during the deployment phase. Both $\mathcal{E}$ and $\mathcal{W}$ are models deployable on the edge side, thus lacking general knowledge about mobile operations. Since $\mathcal{E}$ has already compensated for the lack of general knowledge through few-shot learning with RAG, only $\mathcal{W}$ requires supervised fine-tuning to unleash its potential in query rewriting.
To warm up the trainable scheme query writer $\mathcal{W}$, we employ a data distillation method, allowing $\mathcal{W}$ to learn the query and SOP rewriting logic from large-scale models with strong general knowledge (e.g., GPT-4o). For the support dataset of MobileIAR, we used GPT-4o to replace the query writer $\mathcal{W}$ during the deployment phase, obtaining personalized rewritten queries $\hat{q}$ and personalized rewritten SOPs $\hat{p}$ for each query. Then, using these $\hat{q}$ and $\hat{p}$ as labels, we performed supervised fine-tuning on Qwen3-4B to warm up the query writer $\mathcal{W}$, enabling it to learn how to rewrite queries and SOPs using personalized user information. The training objective can be formulated as:
[…]
After the warm-up training, $\mathcal{W}$ acquires the capability to rewrite queries and SOPs in a user-specific manner.
5. Experiments
5.1. Experiments Setup
5.1.1. Baseline